• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name miham

    Log entry time 00:03:08 on December 4,2009

    Entry number 304497

    keyword=Swing shift Summary (12/03/09)


    Swing shift summary: Miha Mihovilovic (SL), John J. LeRose (TO)
    Simon Sirca (3rd)
    ===================================================================

    EVENTS
    ------
    16:00 The Swing shift started normally. We are taking data.

    17:50 The experiment is running smoothly.

    18:15 MCC called and reported that they have a problem with one
    RF cavity. They were able to bring it up briefly, but now it
    is down again. They have called somebody to fix it, but they
    said that it will probably take a while. They will call us
    back with more details, when the expert arrives.


    18:30 Beam is back. We are taking data again.

    18:51 MCC called and reported that they need to take a beam away
    for approximately 5min, because they need to make some
    adjustments to one of their RF cavities.

    18:57 MCC called and said that they are ready to send beam back.


    21:21 We noticed in the Compton DAQ status window that ROC0 did
    not report status for a while. In spite of that DAQ was
    still counting events. We have decided to do the coda
    reboot and start another run #21387. Now it works
    normally again.

    22:13 We received call from MCC, telling us that they can not see
    our compton and that there might be a problem with it IOC.
    We first tried to reboot coda ROCs but it did not help.
    Therefore we called Brad for help. He noticed that an hour
    ago there mush have been a serious problem with compton. This
    happened exactly when we had to reboot Compton for the first
    time. Brad suggested us to call Alexandre Camsonne. He said
    that he will look into it and that we might have to reboot
    IOC.

    22:40 Alexandre Camsonne came in to fix the compton.

    23:04 Beam HWP was changed from OUT to IN!


    23:53 Compton DAQ seems to be working again. Current
    run number is #21391.







    RUN LIST
    ===================================================================

    Compton DAQ:
    --------------
    #21386: ??:??/21:21, End of run failed! N = 830k
    #21387: 21:21/22:13, End of run failed! N = 470k
    #21391: 23:23/ (Passed to next shift)


    Right Arm DAQ:
    --------------
    5488: 15:58/16:54, I = 100uA, N = 5.7M, DT = ?%, Raster = (4x4)mm
    5489: 16:57/17:54, I = 100uA, N = 5.6M, DT = ?%, Raster = (4x4)mm
    5490: JUNK (Very few events)
    5491: 18:30/19:26, I = 100uA, N = 5.1M, DT = 15%, Raster = (4x4)mm
    5492: 19:28/20:24, I = 100uA, N = 5.7M, DT = 15%, Raster = (4x4)mm
    5493: 20:25/21:54, I = 100uA, N = 5.6M, DT = 15%, Raster = (4x4)mm
    5494: 21:24/22:20, I = 100uA, N = 5.5M, DT = 15%, Raster = (4x4)mm
    5495: 22:22/23:00, I = 100uA, N = 2.5M, DT = 15%, Raster = (4x4)mm
    5496: JUNK due to wrong HV settings for the FeedBack Monitor
    5497: 23:14/??:??, I = 100uA, N = ???M, DT = 15%, Raster = (4x4)mm
    (Run passed to next shift)


    Left Arm DAQ:
    --------------

    25956: 15:58/16:54, I = 100uA, N = 5.4M, DT = 5%, Raster = (4x4)mm
    25957: 16:57/17:54, I = 100uA, N = 5.4M, DT = ?%, Raster = (4x4)mm
    25958: JUNK (Very few events)
    25959: 18:30/19:26, I = 100uA, N = 4.8M, DT = 6%, Raster = (4x4)mm
    25960: 19:28/20:24, I = 100uA, N = 5.4M, DT = 6%, Raster = (4x4)mm
    25961: 20:25/21:54, I = 100uA, N = 5.3M, DT = 6%, Raster = (4x4)mm
    25962: 21:24/22:20, I = 100uA, N = 5.2M, DT = 6%, Raster = (4x4)mm
    25963: 22:22/23:00, I = 100uA, N = 2.3M, DT = 6%, Raster = (4x4)mm
    25964: JUNK due to wrong HV settings for the FeedBack Monitor
    25965: 23:14/??:??, I = 100uA, N = ???M, DT = 7%, Raster = (4x4)mm
    (Run passed to next shift)



    Happex DAQ:
    -----------
    14391: ??:??/16:54, I = 100uA, N = 100k, HWP = OUT
    14392: 16:57/17:53, I = 100uA, N = 100k, HWP = OUT
    14393: 17:56/18:14, I = 31uA, N = 100k, HWP = OUT
    14394: 18:29/19:29, I = 100uA, N = 100k, HWP = OUT
    14395: 19:27/20:24, I = 100uA, N = 100k, HWP = OUT
    14396: 20:25/21:22, I = 100uA, N = 100k, HWP = OUT
    14397: 21:23/22:20, I = 100uA, N = 100k, HWP = OUT
    14398: 22:22/23:00, I = 100uA, N = 68k, HWP = OUT
    14399: 23:05/23:12, I = 100uA, N = 11k, HWP = IN
    14400: 23:14/??:??, I = 100uA, N = ??k, HWP = IN
    (Run passed to next shift)